# Pseudo-Asynchronous Local SGD: Robust and Efficient Data-Parallel Training

### v1

@@ Abstract
**footnotetext: Alphabetical order, –These authors contributed equally to this work. This work was partially done when H.Naganuma and X.Zhang were Microsoft Research interns\$$\dagger$\$$\dagger$\$footnotetext: Alphabetical order
@@ 1 Introduction
  • Empirical Validation: We demonstrate the effectiveness of PALSGD through experiments on ImageNet-1K (Deng et al., 2009), TinyStories (Eldan & Li, 2023), and CIFAR10 datasets. We show that it achieves superior training efficiency compared compared to existing methods like Distributed Data Parallel (DDP) and DiLoCo (Douillard et al., 2023) (Section 7, Figures 1, 2, and 3).
@@ 4.1 Local SGD and Variants
Local SGD is a widely adopted technique for reducing communication overhead in distributed optimization. It allows workers to perform multiple local updates before synchronizing with a central server, minimizing the need for frequent gradient exchanges and improving communication efficiency. Introduced in federated learning by McMahan et al. (2017), Local SGD has been extensively studied for its convergence properties. Stich (2018) demonstrated that it converges at the same rate as mini-batch SGD, particularly in smooth and strongly convex settings. Subsequent works, such as Haddadpour et al. (2019) and Koloskova et al. (2020), extended these results by analyzing Local SGD under more generalized frameworks, including varying network topologies, different convexity settings, and data heterogeneity. The key insight of our algorithm is that it allows workers to mix with the "central model" more frequently without incurring extra communication overhead, leading to better alignment of local models compared to Local SGD.
@@
Several variants of Local SGD have been proposed to further enhance scalability. Elastic Averaging SGD (EASGD) (Zhang et al., 2015) allows local models to diverge from the global model within a bounded range, improving convergence in non-convex settings. While our proposed algorithm similarly introduces slackness through proximal terms in the loss function, it further reduces communication by employing stochastic synchronization rather than stepwise synchronization. Moreover, our algorithm reduces variance between local models more effectively by ensuring that all workers start from the same global model every $H$ local steps, leading to improved consistency across workers. Cooperative SGD (Wang & Joshi, 2021) expands on Local SGD by enabling direct communication between workers, reducing reliance on a central server and improving robustness in decentralized systems. The Scaffnew algorithm (Mishchenko et al., 2022) extends Local SGD by introducing probabilistic communication updates and control variates, with communication handled through a more complex proximal optimization process.
@@
Additionally, Post-Local SGD, a combination of mini-batch SGD and Local SGD, was introduced by Lin et al. (2018) and shown to strike a better balance between communication efficiency and generalization performance in deep learning tasks (Gu et al., 2023). In our proposed method, we adopt a similar strategy to Post-Local SGD by using mini-batch SGD in the warmup phase, followed by the application of our PALSGD algorithm in the second phase. This approach allows us to leverage the fast initial convergence of mini-batch SGD before transitioning to our more communication-efficient method. Unlike FedProx (Li et al., 2020b), which explicitly penalizes local model divergence through a proximal term, our method implicitly controls model alignment via probabilistic pseudo-synchronization, avoiding the need for additional hyperparameter tuning. Furthermore, while our approach shares similarities with L2GD Bergou et al. (2022), we incorporate a warmup phase and an inner-outer loop optimization structure, which are particularly crucial when using AdamW for large-scale vision and language model training.
@@ 5 Proposed Method: PALSGD
Our PALSGD algorithm, outlined in Algorithm 1, extends the Local SGD framework by introducing a pseudo-synchronous step, which allows workers to probabilistically align with a locally stored global model copy instead of requiring immediate full synchronization. Compared to Local SGD, our approach improves training efficiency by reducing synchronization frequency while maintaining model consistency.
5.1 Worker Storage Structure and Local Updates
In our distributed setting, each worker maintains two key model states: a local model $x_{k}^{(t)}$, which is updated using data from its assigned shard, and a global model copy $x^{(t)}$, which stores the last synchronized state and remains unchanged between full synchronization steps. Training progresses over multiple iterations, where each worker performs local updates independently for $H$ steps before synchronizing with the global model.
At each local step, the worker follows one of two possible update mechanisms. With probability $1-p$, it performs a standard gradient-based update using the optimizer InnerOPT, such as SGD or AdamW:
@@ 5.1 Worker Storage Structure and Local Updates
Alternatively, with probability $p$, the worker performs a pseudo-synchronization step, aligning its local model with the stored global model copy using an exponential moving average:
$x_{k}^{(t)}\leftarrow x_{k}^{(t)}-\frac{\alpha_{t}\eta_{t}}{p}(x_{k}^{(t)}-x^{(t)}).$ (5)
This update acts as a lightweight synchronization mechanism, reducing model divergence between workers without requiring communication.
@@ 5.2 Reducing the Frequency of Full Synchronization
Unlike fully synchronous distributed training, where workers communicate at every iteration, PALSGD reduces communication costs by leveraging pseudo-synchronization to partially align workers’ models. This reduces the need for frequent global synchronization and mitigates the problem of idle time, where faster workers must wait for slower ones before proceeding. By allowing workers to continue making local progress, PALSGD improves scalability in large distributed systems while maintaining sufficient alignment to ensure convergence.
After every $H$ iterations, a global synchronization occurs via an All-Reduce operation, which aggregates the model differences across workers:
$\Delta^{(t)}=\textsc{All-Reduce}(x^{(t-1)}-x_{k}^{(t)}).$ (6)
The aggregated differences are then used to update the global model using an outer optimization step:
@@
By performing this synchronization less frequently, PALSGD significantly reduces communication overhead while preserving model consistency.
5.3 Optimizer Strategy and Practical Enhancements
To further improve the empirical performance, we apply several practical techniques. First, similar to Post-Local SGD (Lin et al., 2018), we initialize the model $x^{(0)}$ from a model pretrained by DDP. This addresses the instability issues often observed in the initial phases of training. Additionally, we employ a decoupled optimizer strategy. Following the DiLoCo framework (Douillard et al., 2023), we employed AdamW (Kingma, 2014) as InnerOPT that handles local updates and Nesterov momentum (Sutskever et al., 2013) as OuterOPT that updates the global models. Together, these modifications ensure that PALSGD not only reduces communication costs but also achieves faster convergence and better model performance across various deep learning tasks.
@@ 6 Theoretical Results
$\mathbb{E}[F(\hat{x}_{T})]-F(x^{*})\leq\tilde{O}\Bigg{(}\frac{\sigma^{2}}{\mu KT}+\frac{(\kappa+p^{-1}H)\sigma^{2}}{\mu T^{2}}\Bigg{)}.$ (8)
The convergence bound in Theorem 1 contains two parts. The first term reflects a linear speedup with respect to the number of workers $K$ and is independent of the hyper-parameters for the pseudo-synchronization steps. When $T$ is sufficiently large, this term dominates and matches with the Cramer-Rao bound for estimating a single random variable. The second term arises from the stochastic optimization process. As $K$ increases and communication becomes frequent (i.e., $p^{-1}$ and $H$ are small constants), we can simplify the error bound to $\frac{\kappa\sigma^{2}}{\mu T^{2}}$, which matches the lower bound for stochastic gradient descent in the strongly convex case (Koloskova et al., 2020).
@@ 7.1 Experimental Setup
We performed experiments on two different workloads to assess the performance of PALSGD in a distributed training setup. For image classification, we used ImageNet-1K (Deng et al., 2009) with ResNet-50 (He et al., 2016). For language modeling, we employed TinyStories (Eldan & Li, 2023), using GPT-Neo [^1] https://huggingface.co/docs/transformers/en/model_doc/gpt_neo with 8 million and 125 million parameters. We compared PALSGD with DDP, Local SGD, and DiLoCo across varying numbers of workers and synchronization intervals. For ImageNet-1K experiments, we trained the models with $K=4$ workers, while for the TinyStories experiments, we trained the models using 4 to 8 workers. We also performed additional experiments on CIFAR-10 (Krizhevsky et al., 2009) with a VGG-16 (Simonyan & Zisserman, 2014), including comparisons of training efficiency and accuracy. Details of this experiment are provided in the Appendix C.4 .
To ensure a fair comparison, we evaluated PALSGD alongside three widely used distributed training baselines: Distributed Data Parallel (DDP), Local SGD, and DiLoCo (Douillard et al., 2023). Following the standard approach for post-Local SGD method (Lin et al., 2018), all three algorithms — Local SGD, DiLoCo, and PALSGD - used DDP as warmup during the initial training steps. This warmup phase stabilizes training and improves model initialization before transitioning to a more communication-efficient distributed method. The training hyperparameters of all algorithms, including learning rates and $\eta$, were tuned independently for each choice of $K$ (number of workers) and $H$ (synchronization interval). This tuning process ensures that each algorithm is evaluated under its best-performing configuration for a given distributed training setting. Further details including the ablation studies are provided in Appendix B and  D.
@@ 7.2 Image Classification Tasks: ImageNet-1K on ResNet 50
Figure 1: ImageNet-1K on ResNet-50 Experiments ($K=4$ / $H=64$): PALSGD demonstrates faster training compared to DDP and DiLoCo, while achieving the same target validation accuracy. The results regarding training accuracy are presented in Figure 5 in Appendix C.1.
@@ 7.3 Language Modeling Tasks: TinyStories on GPT-Neo
7.3.1 GPT-Neo-8M Experiments
Figure 2: GPT-Neo-125M Experiments (K=8 / H=16): Training time comparison across distributed algorithm to achieve target validation loss. While PALSGD achieves fastest and lowest loss, DDP is slowest and DiLoCo did not achieve target loss. The results regarding training accuracy are presented in Figure 8 in Appendix C.3.
Figure 3: GPT-Neo-8M Experiments (K=8 / H=16): Training time comparison across distributed algorithm to achieve target validation loss. PALSGD achieves fastest and lowest loss, DDP is slowest and DiLoCo did not achieve target loss. The results regarding training accuracy are presented in Figure 7 in Appendix C.2.2.
@@ 7.3.1 GPT-Neo-8M Experiments
In addition, we performed a series of ablation studies on the hyperparameters of GPT-Neo-8M experiment, and we present the result of synchronization interval $H$ here. Ablation studies on other hyperparameters—including the optimizer, random seeds, $\eta$, and $p$—are provided in Appendix D.
A smaller $H$ leads to longer training times due to more frequent communication during explicit synchronization. Conversely, a larger $H$ reduces the number of All-Reduce operations, thereby shortening the time needed to complete an epoch. To assess the effect of $H$, we conducted an ablation study using GPT-Neo-8M with $K=8$ workers. The results, shown in Figure 4, illustrate how different values of $H$ impact validation loss (left) and training time (right).

### v2

@@ Abstract
[…]
@@ 1 Introduction
  • Empirical Validation: We demonstrate the effectiveness of PALSGD through experiments on ImageNet-1K (Deng et al., 2009), TinyStories (Eldan & Li, 2023), and CIFAR10 datasets. We show that it achieves superior training efficiency compared compared to existing methods like Distributed Data Parallel (DDP) and DiLoCo (Douillard et al., 2023) (Section 7, Figures 2, 3, and 4).
@@ 4.1 Local SGD and Variants
Local SGD is a widely adopted technique for reducing communication overhead in distributed optimization. It allows workers to perform multiple local updates before synchronizing with a central server, minimizing the need for frequent gradient exchanges and improving communication efficiency. Introduced in federated learning by McMahan et al. (2017a), Local SGD has been extensively studied for its convergence properties. Stich (2018) demonstrated that it converges at the same rate as mini-batch SGD, particularly in smooth and strongly convex settings. Subsequent works, such as Haddadpour et al. (2019) and Koloskova et al. (2020), extended these results by analyzing Local SGD under more generalized frameworks, including varying network topologies, different convexity settings, and data heterogeneity. Similarly, Yu et al. (2019) provide an analysis of Local SGD, which they term Parallel Restarted SGD, further demystifying the effectiveness of model averaging in deep learning. The key insight of our algorithm is that it allows workers to mix with the "central model" more frequently without incurring extra communication overhead, leading to better alignment of local models compared to Local SGD.
@@
Several variants of Local SGD have been proposed to further enhance scalability. Elastic Averaging SGD (EASGD) (Zhang et al., 2015) allows local models to diverge from the global model within a bounded range, improving convergence in non-convex settings. While our proposed algorithm similarly introduces slackness through proximal terms in the loss function, it further reduces communication by employing stochastic synchronization rather than stepwise synchronization. Moreover, our algorithm reduces variance between local models more effectively by ensuring that all workers start from the same global model every $H$ local steps, leading to improved consistency across workers. Cooperative SGD (Wang & Joshi, 2021) expands on Local SGD by enabling direct communication between workers, reducing reliance on a central server and improving robustness in decentralized systems. The Scaffnew algorithm (Mishchenko et al., 2022) extends Local SGD by introducing probabilistic communication updates and control variates, with communication handled through a more complex proximal optimization process. However, its reliance on exact control-variate updates makes it difficult to extend to modern optimizers like AdamW without significant modifications to the algorithm and its memory footprint. In contrast, our approach uses a simpler control mechanism, making PALSGD substantially easier to integrate into existing deep learning pipelines.
@@
Additionally, Post-Local SGD, a combination of mini-batch SGD and Local SGD, was introduced by Lin et al. (2018) and shown to strike a better balance between communication efficiency and generalization performance in deep learning tasks (Gu et al., 2023). In our proposed method, we adopt a similar strategy to Post-Local SGD by using mini-batch SGD in the warmup phase, followed by the application of our PALSGD algorithm in the second phase. This approach allows us to leverage the fast initial convergence of mini-batch SGD before transitioning to our more communication-efficient method. SCAFFOLD (Karimireddy et al., 2020), addresses the issue of client drift by using control variates to correct the local update direction, which differs from our probabilistic mixing approach. In addition, SCAFFOLD is primarily designed to address non-IID scenarios and has been reported to perform worse than FedAvg (McMahan et al., 2017b) in IID settings. Another key related work, FedProx (Li et al., 2020b), explicitly penalizes local model divergence through a proximal term. Unlike the above-mentioned works, our method implicitly controls model alignment via probabilistic pseudo-synchronization, avoiding the need for additional hyperparameter tuning and extra forward-backward calculation. Our method also contrast with synchronous approaches like DropCompute (Giladi et al., 2023), which handles stragglers by dropping slow gradient computations rather than extending the local update period Furthermore, while our approach shares similarities with L2GD Bergou et al. (2022), we incorporate a warmup phase and an inner-outer loop optimization structure, which are particularly crucial when using AdamW for large-scale vision and language model training.
@@ 4.4 Asynchronous and Pseudo-Asynchronous Methods
Recent works have also explored fully asynchronous and decentralized settings to further mitigate synchronization bottlenecks. Methods like Shadowheart SGD (Tyurin et al., 2024) and SWIFT (Bornstein et al., 2022) leverage parameter-server architectures to enable wait-free communication, while others have explored techniques like model fragmentation to boost asynchronous learning (Biswas et al., 2025). These approaches differ from our semi-synchronous framework, which maintains periodic global synchronization points to avoid issues like unbounded gradient staleness.
@@ 5 Proposed Method: PALSGD
The primary challenge in extending the synchronization interval $H$ in Local SGD is the problem of model divergence, where worker models drift from the global optimum and degrade convergence. To address this, we propose Pseudo-Asynchronous Local SGD (PALSGD), a novel extension of the Local SGD framework. PALSGD introduces a pseudo-synchronous step that acts as a lightweight, communication-free regularizer to explicitly counteract this divergence. As detailed in Algorithm 1, this step allows workers to probabilistically align with a locally stored global model copy, enabling the use of longer synchronization intervals without sacrificing model consistency.
5.1 Probablistic Regularization via Pseudo-Synchronization
In our distributed setting, each worker maintains two key model states: a local model $x_{k}^{(t)}$, which is updated using data from its assigned shard, and a global model copy $x^{(t)}$, which stores the last synchronized state and remains unchanged between full synchronization steps. Training progresses over multiple iterations, where each worker performs local updates independently for $H$ steps before a full synchronization occurs.
At each local step, the worker follows one of two paths based on a probability $p$. With probability 1p1-p, it performs a standard gradient-based update using its local data and the optimizer InnerOPT:
@@ 5.1 Probablistic Regularization via Pseudo-Synchronization
Alternatively, with probability $p$, the worker performs a pseudo-synchronization update:
$x_{k}^{(t+1)}\leftarrow x_{k}^{(t)}-\frac{\alpha_{t}\eta_{t}}{p}(x_{k}^{(t)}-x^{(t)}).$ (5)
This update is the core of PALSGD and serves as a principled mechanism to control model divergence. Its form can be justified in several ways:
  • As a Proximal Regularizer: The update in Equation (2) can be interpreted as a single gradient step on a quadratic penalty term, $\frac{\eta_{t}}{2}\|x_{k}-x^{(t)}\|^{2}$. This term explicitly regularizes the local model by pulling it towards the last-known global state $x^{(t)}$. It penalizes deviation from this shared reference point, thereby constraining the local optimization paths and preventing workers from drifting too far apart.
  • As an Exponential Moving Average (EMA): The update can be rewritten as a mixing step: $x_{k}^{(t+1)}\leftarrow(1-\beta)x_{k}^{(t)}+\beta x^{(t)}$, where the mixing coefficient is $\beta=\frac{\alpha_{t}\eta_{t}}{p}$. This shows that the worker’s parameters are an exponential moving average of their own trajectory and the global trajectory. This form is well-known for stabilizing updates by dampening oscillations and reducing the variance of the parameter updates.
By probabilistically applying this step, PALSGD ensures that local models remain sufficiently aligned, acting as a lightweight mechanism to reduce model divergence without incurring any communication cost. We refer to this as a pseudo-synchronization step to distinguish it from a true synchronization event like All-Reduce. In a true synchronization, workers exchange information and compute a global average, which requires all workers to pause and communicate. In contrast, a pseudo-synchronization step is a purely local operation. Each worker moves its local model towards its own stored copy of the global model, which may be stale by up to (H1)(H-1) steps. This step involves no communication with other workers, thus avoiding network latency and idle time for faster workers.
@@ 5.2 Reducing the Frequency of Full Synchronization
By employing the probabilistic regularization described above to maintain model consistency, PALSGD can safely reduce the frequency of full, communication-heavy synchronizations. Unlike fully synchronous methods where workers communicate at every iteration, PALSGD leverages pseudo-synchronization to partially align workers’ models, reducing the need for frequent global updates. This approach mitigates the problem of idle time, where faster workers must wait for slower ones, and improves scalability, especially in large distributed systems.
A full global synchronization occurs only every $H$ iterations via an AllReduce operation, which aggregates the model differences across all $K$ workers:
$\Delta^{(t)}=\textsc{AllReduce}(x^{(t-1)}-x_{k}^{(t)}).$ (6)
The aggregated difference $\Delta^{(t)}$ is then used to update the global model with an outer optimization step:
@@
By performing this costly operation less frequently, PALSGD significantly reduces communication overhead while the pseudo-synchronization mechanism preserves model consistency between these updates.
5.3 Practical Enhancements and Implementation Details
To further improve empirical performance and ensure stability, PALSGD incorporates several practical techniques with minimal overhead. First, similar to Post-Local SGD (Lin et al., 2018), we initialize training with a DDP warm-up phase to address potential instability when gradients are large and noisy. Following this, we employ a decoupled optimizer strategy, as validated in the DiLoCo framework (Douillard et al., 2023). We use AdamW (Kingma, 2014) as the InnerOPT for rapid local progress and Nesterov momentum (Sutskever et al., 2013) as the OuterOPT to robustly aggregate updates.
From an implementation standpoint, these enhancements are lightweight:
  • Memory Requirements: Each worker stores the local model $x_{k}^{(t)}$, the global model copy $x^{(t)}$, and the outer optimizer state. This overhead, equivalent to one extra model and its optimizer state, is well within the capacity of modern GPUs.
  • Communication Requirements: The pseudo-synchronization step is a purely local computation and adds no communication overhead. Communication occurs only during the All-Reduce operation every $H$ steps. The primary efficiency gain of PALSGD stems from its ability to use a larger $H$ effectively, thereby reducing the total number of these expensive communication events.
Together, these modifications ensure that PALSGD not only reduces communication costs but also achieves faster convergence and better model performance across various deep learning tasks.
@@ 6 Theoretical Results
$\mathbb{E}[F(\hat{x}_{T})]-F(x^{*})\leq\tilde{O}\Bigg{(}\frac{\sigma^{2}}{\mu KT}+\frac{\kappa H^{2}\sigma^{2}}{\mu T^{2}}\Bigg{)}.$ (8)
The convergence bound in Theorem 1 contains two parts. The first term, $\frac{\sigma^{2}}{\mu KT}$, reflects a linear speedup with respect to the number of workers $K$ and is independent of the hyperparameters for the pseudo-synchronization steps. When $T$ is sufficiently large, this term dominates and matches the Cramer-Rao bound for estimating a single random variable. The second term arises from the stochastic optimization process. As $K$ increases and communication becomes frequent (i.e., $H$ is small constant), we can simplify the error bound to $\frac{\kappa\sigma^{2}}{\mu T^{2}}$, which matches the lower bound for stochastic gradient descent in the strongly convex case (Koloskova et al., 2020).
@@ 7.1 Experimental Setup
We performed experiments on five different workloads to assess the performance of PALSGD in a distributed training setup. For image classification, we used CIFAR-10 (Krizhevsky et al., 2009) with small CNN and VGG16(Simonyan & Zisserman, 2014), ImageNet-1K (Deng et al., 2009) with ResNet-50 (He et al., 2016). For language modeling, we employed TinyStories (Eldan & Li, 2023), using GPT-Neo [^1] https://huggingface.co/docs/transformers/en/model_doc/gpt_neo with 8 million and 125 million parameters.
To ensure a fair comparison, we evaluated PALSGD alongside three widely used distributed training baselines: Distributed Data Parallel (DDP), Local SGD, and DiLoCo (Douillard et al., 2023). In the CIFAR-10 experiments, we varied the number of workers from 4 to 64. For ImageNet-1K experiments, we trained the models with $K=4$ workers, while for the TinyStories experiments, we trained the models using 4 to 8 workers.
Following the standard approach for post-Local SGD method (Lin et al., 2018), all three algorithms — Local SGD, DiLoCo, and PALSGD - used DDP as warmup during the initial training steps. This warmup phase stabilizes training and improves model initialization before transitioning to a more communication-efficient distributed method. The training hyperparameters of all algorithms, including learning rates and $\eta$, were tuned independently for each choice of $K$ (number of workers) and $H$ (synchronization interval). This tuning process ensures that each algorithm is evaluated under its best-performing configuration for a given distributed training setting. Further details including the ablation studies are provided in Appendix B and  E.
7.2 Preliminary Experiments: CIFAR10 on Small CNN
We begin by presenting small-scale results, including comparisons with Local SGD (Ortiz et al., 2021). DiLoCo (Douillard et al., 2023) can be interpreted as equivalent to Local SGD, or to FedAvg (McMahan et al., 2017b) in the IID case, when the outer optimizer is set to SGD. As shown in Figure 6 of the Douillard et al. (2023), DiLoCo outperforms these methods. We conducted preliminary experiments to examine how accuracy changes as the number of workers increases for DDP, Local SGD, DiLoCo, and PALSGD. The results are shown in Figure 1. The synchronization interval was fixed at 32.
The results indicate that LocalSGD degrades substantially as the number of workers increases, compared to the other three methods. Our result consistents with observations in Lin et al. (2018). DiLoCo and PALSGD show similar performance when worker counts or synchronization intervals are small. In contrast, PALSGD maintains better performance when synchronization intervals become large. This occurs because DiLoCo permits extended intervals without ensuring model consistency. PALSGD addresses this issue by imposing pseudo-synchronization constraints.
Based on these findings, we do not include Local SGD in large-scale experiments. Instead, we use DDP as the baseline and DiLoCo as the main comparison method for evaluating PALSGD. Methods such as SCAFFOLD(Karimireddy et al., 2020) are excluded, since prior work has shown that they underperform LocalSGD in the IID setting with reasonable epochs budget (see Table 3 in Karimireddy et al. (2020)).
The following subsections report results from more practical experiments on ImageNet and TinyStories.
Figure 1: Comparison of number of workers across algorithm / simulation experiments with CIFAR10 dataset
@@ 7.3 Image Classification Tasks: ImageNet-1K on ResNet 50
Figure 2: ImageNet-1K on ResNet-50 Experiments ($K=4$ / $H=64$): PALSGD demonstrates faster training compared to DDP and DiLoCo, while achieving the same target validation accuracy. The results regarding training accuracy are presented in Figure C.1 in Appendix C.1.
@@ 7.4 Language Modeling Tasks: TinyStories on GPT-Neo
Figure 3: GPT-Neo-125M Experiments (K=8 / H=16): Training time comparison across distributed algorithm to achieve target validation loss. While PALSGD achieves fastest and lowest loss, DDP is slowest and DiLoCo did not achieve target loss. The results regarding training accuracy are presented in Figure C.4 in Appendix C.3.
Figure 4: GPT-Neo-8M Experiments (K=8 / H=16): Training time comparison across distributed algorithm to achieve target validation loss. PALSGD achieves fastest and lowest loss, DDP is slowest and DiLoCo did not achieve target loss. The results regarding training accuracy are presented in Figure C.3 in Appendix C.2.2.
@@
In addition, we performed a series of ablation studies on the hyperparameters of GPT-Neo-8M experiment, and we present the result of synchronization interval $H$ here. Ablation studies on other hyperparameters—including the optimizer, random seeds, $\eta$, and pp—are provided in Appendix E. Overall, across all workloads we observed a consistent trend: values of $p\geq 0.25$ led to faster training but degraded performance, while in practical applications such as ImageNet-1K and TinyStories, values around 0.05 yielded better results.
For synchronization interval, a smaller $H$ leads to longer training times due to more frequent communication during explicit synchronization. Conversely, a larger $H$ reduces the number of All-Reduce operations, thereby shortening the time needed to complete an epoch. To assess the effect of $H$, we conducted an ablation study using GPT-Neo-8M with $K=8$ workers. The results, shown in Figure 5, illustrate how different values of $H$ impact validation loss (left) and training time (right).